# Finding Dino: A plug-and-play framework for unsupervised detection of out-of-distribution objects using prototypes

### v1

@@ 4.2 Experimental Setup
PROWL is a zero-shot method, which means it does not rely on any training but merely some inference steps. The ViT-S/14 of DINO-v2[27] pre-trained on a dataset of 142 million images is taken as the feature extractor of PROWL. We observe that with a default inverse cosine similarity threshold (Eq. 4) at and CutLER detector threshold of , we obtain a general satisfactory performance across all the datasets. We further report that depending upon the size of the OOD objects and complexity of the OOD datasets, this optimal threshold tends to suffice even at higher values as shown in the ablation study for OOD datasets for Cityscapes (Sec 5.3).
@@ 4.3 Evaluation metric
To the best of our knowledge, there exists no prior work on unsupervised OOD detection from multiple objects in a scene. Thus, there are no established benchmarks and evaluation metrics for the task. The Area under the Receiver Operating Characteristic (AUROC) curve is a standard metric for OOD detection for image classification. However, in the case of multiple OOD object detection and segmentation tasks with a highly imbalanced number of OOD objects, the Area under the Precision-Recall Curve (AUPRC) is a more suitable choice. Thus, to compare with Anomaly Segmentation metrics provided with the SMIYC benchmark, we use AUPRC and False Positive Rate at True Positive Rate of (FPR) metrics for pixel-wise evaluation. Further, for evaluation as a binary segmentation task, we select a fixed inverse cosine similarity threshold for all the datasets and predict (positive) for pixels determined as OOD otherwise . This binary prediction mask is compared with the GT mask and evaluated using mean Intersection over Union (mIoU) and F1 score.
@@ 5 Results and Discussion
In this section, we show different experiments to evaluate the performance of PROWL for OOD detection. In Sec. 5.1, we present results comparing with State-of-the-Art (SOTA) supervised methods in the road driving scene reported in the existing anomaly segmentation benchmarks SMIYC [7]. In Sec. 5.2, we show the applicability of PROWL into other domains such as rail and maritime scenes where there are no directly comparable benchmarks. Lastly, in Sec 5.3, we provide ablation for different hyper-parameters and strategies used in our implementation.
We report results for our method PROWL in two scenarios as explained in Sec. 3.3: (a) PROWL (Without foreground masks) - OOD pixels are directly obtained by thresholding the scores from prototype heatmaps, (b) PROWL + refinement module using foreground masks - The foreground masks generated by unsupervised segmentation methods (STEGO [18], CutLER [31] as applicable) are combined with prototype heatmaps, such that individual masks (not pixels) are detected as OOD objects or not.
5.1 Comparison with the State-of-the-Art on road driving scene
In the road driving scene, we evaluate on datasets with real OOD objects like RoadAnomaly21 andRoadObstacle21 provided by SMIYC benchmark [7] and the synthetic OOD objects in FS-Static given in Fishyscapes [3] benchmark. We created prototype bank for road scene using ODD classes from Cityscapes, in accordance with the classes recommended in SMIYC following [10].
RoadAnomaly | FS Static | RoadObstacle | ||||||
Method | Seg. type | Aux. data | AUPR | FPR | AUPR | FPR | AUPR | FPR |
Supervised | ||||||||
PEBAL | Pixel | ✓ | 45.1 | 44.6 | 92.1 | 1.5 | - | - |
Max. Entropy | Pixel | ✓ | 79.7 | 19.3 | 76.3 | 7.1 | - | - |
DenseHybrid | Pixel | ✓ | 63.9 | 43.2 | 60.0 | 4.9 | - | - |
M2A | Mask | ✓ | 79.7 | 13.5 | - | - | - | - |
@@ 5.1 Comparison with the State-of-the-Art on road driving scene
PROWL + CutLER | Mask | ✗ | 75.25 | 1.75 | 73.29 | 1.46 | 75.43 | 1.84 |
To the best of our knowledge, there are no other unsupervised OOD object detection methods comparable as a baseline to our method. Thus, it should be noted that all the methods used for comparison with SOTA are supervised, i.e. these models are trained on semantic annotations of the classes of Cityscapes, whereas we rely on few inference steps based on prototype features from DINOv2.
@@
In Table 1, we compare results reported by other supervised methods on the validation set of respective OOD datasets. The metrics used are threshold-independent metrics such as AUPR and FPR at the best TPR (or ). We note that while the other reported metrics are based on probabilistic confidence scores, we calculate positives (OOD pixels) by varying the inverse cosine similarity threshold in Eq. 4. Thus, we are unable to compare with the test set of the benchmark due to non-compatibility with expected probabilistic inputs on the official benchmark code. We indicate the segmentation type of each method as pixel or mask-based. Further, many SOTA methods achieve superior results by discriminatively training with auxiliary data, i.e., exposed to few OOD samples as negative data. We include those SOTA methods with backbones having similar predictive power in evaluation and are ranked higher in the SMIYC benchmark. For comparison purposes, we assume Maskomaly [1] as our direct supervised baseline as it is also a zero-shot inference based approach like PROWL, depending on pre-trained Mask2Former [9] model on Cityscapes. Unlike other methods, it provided SOTA results on RoadAnomaly amongst the methods not trained with auxiliary data.
We show that PROWL with CutLER performs best compared to all the other variants amongst the unsupervised methods. In comparison to supervised methods trained without auxiliary data, PROWL outperforms on the RoadAnomaly and RoadObstacle datasets. It provides overall best FPR for RoadAnomaly across all methods and gain in AUPR. For FS Static, our unsupervised methods perform quite comparable to supervised methods (second best) and performs better than Maskomaly as well. We note that most supervised methods do not report results on validation set for RoadObstacles and the reproduced results for Maskomaly show poor performance, also shown in Fig 3.
RoadAnomaly | FS Static | RoadObstacle | |||||
Method | Supervised | mIoU | F1 | mIoU | F1 | mIoU | F1 |
Maskomaly* | Yes | 46.86 | 56.76 | 23.97 | 29.92 | 9.31 | 11.29 |
PROWL | No | 38.46 | 54.68 | 39.43 | 50.75 | 6.79 | 11.21 |
PROWL + STEGO | No | 56.24 | 69.81 | 39.21 | 49.9 | 29.92 | 35.10 |
PROWL + CutLER | No | 75.22 | 85.25 | 64.79 | 72.18 | 49.16 | 53.31 |
@@
In Table 2 and Fig. 3, we present the comparative performance of our methods against supervised SOTA baseline (Maskomaly [1]) as a binary segmentation task. For this, we compare the binary GT masks with OOD pixels given as with the binary masks generated from the pixels predicted as OOD by different methods. We evaluate mIoU and F1 scores with fixed inverse cosine similarity thresholds on the best threshold for validation set of respective OOD datasets for Cityscapes (Fig. 4(c)). The authors of Maskomaly [1] report results on the validation set by calculating best threshold per-image by checking best F1 score using GT. However, for fair comparison we reproduce all the results for Maskomaly with their reported fixed confidence threshold of .
@@
In Table 2, we show that PROWL with CutLER outperforms other unsupervised variants as well as supervised baseline in terms of both mIoU and F1 scores with a significant margin. We observe that there’s an overall decreasing trend of performance in most methods across OOD datasets from RoadAnomaly, FS-Static to RoadObstacle. This can be attributed also to the difficulty of the dataset and the size of OOD object within the image. We can correlate this with the qualitative results provided in Fig. 3. RoadAnomaly dataset provides real OOD objects which are comparably big like the carriage or helicopter but also in a scene different than the city roads in Cityscapes. Thus, while bigger OOD objects are easier to detect, the background objects might also be often detected as OOD. We observe this trend for results corresponding to prototype heatmaps in PROWL and STEGO with PROWL. Although Maskomaly reports good results on SMIYC benchmark, we still observe that the masks detected for OOD are highly insufficient with only partial detections. In contrast, since CutLER detects foreground object boxes and then provides semantic masks, the entire mask is precisely detected as OOD object using PROWL. For FS-Static, the OOD objects are synthetic images which often have a matching texture with the background of Cityscapes test images. So, although most methods do not falsely predict too many background pixels as OOD, we see some false predictions in the background for Maskomaly and PROWL in the image of dog crossing the road. The primary difficulty in this dataset could be missing the OOD object due to similar texture as background. This is apparent for the Maskomaly image with dog sitting on road. Overall, STEGO with PROWL performs quite nicely, as STEGO has learnt via contrastive clustering to find semantic correspondences in foreground masks on Cityscapes and thus nicely finds small semantic masks which are significantly different rather than bigger deviations in the scene as in RoadAnomaly. We observe that CutLER with PROWL precisely detects the OOD masks in most cases. For the RoadObstacle dataset, the OOD objects are meant as real obstacles of small sizes and varying numbers placed on the ego-track at different distances. Due to smaller sizes with increasing distance, all methods show worse performance as compared to other datasets. The reproduced Maskomaly results falsely captures most of the road. PROWL finds the OOD objects in most cases but also falsely predicts many background pixels as OOD. We observe the superior qualitative performance of object masks provided by STEGO and CutLER which helps in localising the objects as entire masks and then detecting them as OOD via PROWL. We note although STEGO localises all the OOD object masks, the semantic masks are somewhat discontinuous. In comparison, CutLER with PROWL precisely localises and segments the OOD object masks, justifying the overall better performance. Although vanilla PROWL is quite close to detecting the OOD pixels, however with above examples we show that the additional unsupervised foreground masks helping in refining the OOD object masks avoiding spurious false predictions.
5.2 Evaluation of plug and play application to other domains
To demonstrate zero-shot plug-and-play application of PROWL to other domains, we extend our evaluation to rail and maritime ODD scene with respective OOD datasets given in Sec. 4.1. For rail ODD, we use the dataset created with inpainted OOD objects. For simplicity, we consider the following predominant classes in RailSem19 [33] for creating prototype bank of known classes - train car, platform, rail, fence, person, pole. Similarly, for maritime we refer to MODD [4] and create prototype bank with sky, sea as known ODD classes. Every other object is OOD. Since there exists no benchmark on RailSem19 OOD detection and the MODS [5] uses separate evaluation strategy, we provide evaluation using our zero-shot methods with PROWL. Since CutLER models were trained without labels only on ImageNet [11], it can easily provide zero-shot inference on any domain, whereas STEGO still needs to train on the datasets of new domain although without labels. Similarly, our supervised baseline Maskomaly [1] relies on Mask2Former which still needs to train with the labels of the respective training data, thus can not be readily used for comparison.
@@ 5.2 Evaluation of plug and play application to other domains
Method | Supervised | mIoU | F1 | mIoU | F1 |
PROWL | No | 8.01 | 14.33 | 62.35 | 76.36 |
PROWL + CutLER | No | 83.29 | 90.38 | 73.30 | 84.08 |
@@
In Table 3, we show the performance of PROWL vs CutLER with PROWL as a binary segmentation task using fixed similarity threshold for detection of OOD pixels. We find that CutLER with PROWL performs better as compared to vanilla PROWL in both cases. In Fig. 4, we show corresponding qualitative results. We observe that PROWL is significantly worse in case Rail Inpainted OOD data as compared to MODD where it performs quite comparably as CutLER with PROWL. This could be attributed to the fact that prototype heatmaps in PROWL provide per-pixel outputs, causing false positives to the background objects like vegetation, buildings that are not in the assumed ODD class list. CutLER provides an advantage here, where the relevant object masks (including OOD) can be filtered based on foreground score while irrelevant background objects can be ignored. However, in MODD dataset, the task is relatively simple as everything other than sea, sky should be detected as OOD or obstacle, thus both methods perform well on this dataset.
5.3 Ablations
In Fig.5, we show ablation study for different hyperparameters used in the experiments with PROWL depending upon complexity of different OOD datasets, such as variation of average OOD object size in pixels Fig. 4(a) where RoadAnomaly shows much bigger objects as compared to RoadObstacle with smaller objects. Fig.4(b) shows CutLER thresholds for generating optimal number of foreground masks for PROWL with the inverse cosine similarity threshold set at . Optimal mIoU is achieved for CutLER thresholds of , , and for RoadAnomaly, FS Static, and RoadObstacle datasets respectively. This trend is justifiable as smaller objects require lesser thresholds so as to detect maximum object masks, whereas bigger objects can be reliably detected even with higher detector thresholds. In Fig. 4(c), we fix the CutLER thresholds given above and find the optimum inverse cosine similarity threshold for best OOD detection based on mIoU. We observe FS Static peaks at relatively high threshold of 0.7 as the dataset is quite similar to ID Cityscapes data, whereas RoadAnomaly and RoadObstacle dataset peak at , respectively close to the default threshold at .
Lastly Fig. 6, shows the performance of PROWL with varying number of prototype samples used to create the prototype feature bank for the Cityscapes dataset in road driving domain. We show already quite good AUPRC is achieved with as less as prototypes and it starts to saturate from and above. For our experiments, samples per ODD class were used for Cityscapes. For RailSem19 and MODD datasets with less complexity and lesser number of ODD classes, approximately prototype samples were found to be sufficient.

### v2

@@ 4.2 Experimental Setup
PROWL is completely based on inference on frozen features from foundation models without using any domain data for training.The chosen feature extractor is ViT-S/14 of DINO-v2[28] pre-trained on a dataset of 142 million images. Using a held-out validation set (unused samples from train split), we observed good performance with a default INCS threshold (Eq. 5) of and a low CutLER detector threshold of . We note that for zero-shot OOD detection keeping a low detector thresholds helps in detecting all possible sizes of object instances. We further report that depending upon the size of the OOD objects and complexity of the OOD datasets, this threshold can be further optimised to suffice even at higher values as shown in the ablation study (Sec. 5.3).
@@ 4.3 Evaluation metric
To the best of our knowledge, there exists no prior work, benchmarks and evaluation metrics on zero-shot inference on foundational models for detection of OOD objects. Hence, we compare with existing road anomaly segmentation benchmarks which provide OOD datasets and metrics, such as AUPRC and False Positive Rate at True Positive Rate of (FPR) metrics for pixel-wise evaluation. Further, for evaluation as a binary segmentation task, we select a fixed INCS threshold for all the datasets and predict (positive) for pixels determined as OOD otherwise . This binary prediction mask is compared with the GT mask and evaluated using Intersection over Union (IoU) for the OOD class and F1 score.
@@ 5 Results and Discussion
In this section, different experiments to evaluate the performance of PROWL for OOD detection are shown. In Sec. 5.1, we present results comparing with State-of-the-Art (SOTA) supervised methods in the road driving scene. In Sec. 5.2, we show the applicability of PROWL into other domains such as rail and maritime scenes where there are no directly comparable benchmarks. Lastly, in Sec 5.3, we provide ablation for different hyper-parameters and strategies used in our implementation.
We report results for our method PROWL in two scenarios: (a) PROWL (Without foreground masks in Sec. 3.2) - OOD pixels are directly obtained by thresholding the INCS scores from prototype heatmaps, (b) PROWL + refinement module using foreground masks (Sec. 3.3) - The foreground masks generated by unsupervised segmentation methods (STEGO [19], CutLER [33] as applicable) are combined with prototype heatmaps, such that individual masks (not pixels) are detected as OOD objects or not.
| RoadAnomaly | FS Static | RoadObstacle | ||||||||
| Method | Seg. type | Aux. data | AUPR | FPR | AUPR | FPR | AUPR | FPR | ||
| Supervised | ||||||||||
| PEBAL | Pixel | ✓ | 45.1 | 44.6 | 92.1 | 1.5 | - | |||
| Max. Entropy | Pixel | ✓ | 79.7 | 19.3 | 76.3 | 7.1 | - | - | ||
| DenseHybrid | Pixel | ✓ | 63.9 | 43.2 | 60.0 | 4.9 | - | - | ||
| M2A | Mask | ✓ | 79.7 | 13.5 | - | - | - | - | ||
| RbA | Mask | ✓ | 85.42 | 6.92 | ||||||
@@
[…]
| PROWL + CutLER | Mask | ✗ | 75.25 | 1.75 | 70.27 | 8.21 | 73.53 | 5.58 | ||
5.1 Comparison with SOTA on road driving scene
In the road driving scene, we evaluate on datasets with real OOD objects like RA and RO provided by SMIYC [7] and the synthetic OOD objects in FS-Static given in Fishyscapes [3] benchmark. We created prototype bank for road scene using ODD classes from Cityscapes[10] and also followed in SMIYC.
There exists no prior work serving as a baseline for zero-shot OOD object detection task based on foundation models, it should be noted that all the methods used for comparison with SOTA are fully supervised, i.e. these models are trained with GT instance masks of the classes of Cityscapes, whereas we rely on few inference steps on frozen self-supervised DINOv2 models. We note benchmark (SMIYC, Fishyscapes) evaluation code requires logits based on finite number of closed set categories by varying detector thresholds. However, PROWL is based on zero-shot inference on foundation models which can not output logits on fixed classes making zero-shot methods not directly comparable with benchmarks designed for supervised methods. To still compare with SOTA methods, we calculate evaluation metrics using INCS score (Eq. 5) and regard the corresponding threshold as the variable parameter which is different from detector thresholds. Since SMIYC test GT are not available without benchmark submissions, we report quantitative results only on RA and RO (where GT given) in Table 1, 2 and visually impressive qualitative results on SMIYC test set in Fig. 1, and Fig. 4, 5 in Supplementary.
@@ 5.1 Comparison with SOTA on road driving scene
In Table 1, we compare with SOTA methods using threshold-independent evaluation metrics such as AUPR and FPR at the best TPR (or ). We include those highly-ranked SOTA methods from SMIYC benchmark which also report results on mentioned datasets. We indicate the segmentation type of each method as pixel or mask-based. Further, many SOTA methods achieve superior results by discriminative training with select OOD samples (auxiliary data) during training or exposing to such OOD samples for fine-tuning. For comparison purposes, we assume Maskomaly [1] as our direct supervised baseline as it is also a zero-shot inference based approach like PROWL, depending on pre-trained Mask2Former [9] on Cityscapes, provides reproducible code for considered datasets and does not depend on auxiliary OOD data. We show that PROWL with CutLER performs best compared to all the other variants amongst the zero-shot methods. In comparison to supervised methods trained without auxiliary data, PROWL with CutLER provides overall best FPR and second best to cDNP[15] in AUPR on the RA and overall outperforms on RO datasets. We note that most supervised methods do not report results on RO and the reproduced results for Maskomaly show poor performance, also shown in Fig. 3. For FS Static, our zero-shot methods performs better than Maskomaly and comparably similar to other supervised methods.
| RoadAnomaly | FS Static | RoadObstacle | ||||
| Method | IoU | F1 | IoU | F1 | IoU | F1 |
| Maskomaly* | 46.86 | 56.76 | 23.97 | 29.92 | 9.31 | 11.29 |
| PROWL | 38.46 | 54.68 | 39.43 | 50.75 | 6.79 | 11.21 |
| PROWL + STEGO | 56.24 | 69.81 | 39.21 | 49.9 | 29.92 | 35.10 |
| PROWL + CutLER | 75.22 | 85.25 | 64.79 | 72.18 | 49.16 | 53.31 |
@@
In Table 2 and Fig. 3, we present the comparative performance of our methods against supervised SOTA baseline (Maskomaly [1]) as a binary segmentation task. For this, the binary GT masks (where OOD pixels are given as ) are compared with the binary masks generated from the predicted OOD pixels via different methods. We evaluate mIoU and F1 scores with fixed INCS thresholds. The authors of Maskomaly [1] report results on the validation set by calculating best threshold every image by checking best F1 score using GT. However, for fair comparison we reproduce all the results for Maskomaly with their reported best fixed confidence threshold of for those datasets.
@@
In Table 2, we show that PROWL with CutLER outperforms other zero-shot variants as well as supervised baseline in terms of both mIoU and F1 scores with a significant margin. We observe that there’s an overall decreasing trend of performance in most methods across OOD datasets from RA, FS-Static to RO. This can be attributed also to the difficulty of the dataset and the size of OOD object within the image. We can correlate this with the qualitative results provided in Fig. 3. RA dataset provides real OOD objects which are comparably big like the carriage or helicopter but also in a scene different than the city roads in Cityscapes. Thus, while bigger OOD objects are easier to detect, the background objects might also be often detected as OOD. We observe this trend for results corresponding to prototype heatmaps in PROWL with STEGO. Although Maskomaly reports good results on SMIYC benchmark, we still observe that the masks detected for OOD are highly insufficient with only partial detections. In contrast, since CutLER first detects foreground object boxes and then provides semantic masks, the entire mask is precisely detected as OOD object using PROWL. For FS-Static, the OOD objects are synthetic images which often have a matching texture with the background of Cityscapes images. Thus, we see some false predictions with the background pixels for Maskomaly and PROWL in the image of dog crossing the road. The primary difficulty in this dataset could be missing the OOD object due to similar texture as background. This is apparent for the Maskomaly image with dog sitting on road. We observe that PROWL with CutLER most precisely detects the OOD masks in most cases. For the RO dataset, the OOD objects are meant as real obstacles of small sizes and varying numbers placed on the ego-track at different distances. Due to smaller sizes with increasing distance, all methods show worse performance as compared to other datasets. The reproduced Maskomaly results falsely captures parts of the road. PROWL finds the OOD objects in most cases but also falsely predicts many background pixels as OOD. We observe the superior qualitative performance of object masks provided by STEGO and CutLER which helps in localising the objects as entire masks and then detecting them as OOD via prototype heatmaps in PROWL. We note although STEGO localises all the OOD object masks, the semantic masks are somewhat discontinuous. In comparison, PROWL with CutLER precisely localises and segments the OOD object masks, justifying the overall better performance. Although vanilla PROWL is quite close to detecting the OOD pixels, however with above examples we show that the additional unsupervised foreground masks helping in refining the OOD object masks avoiding spurious false predictions.
To show generalisability to totally different domain and OOD objects, we further show impressive results on one of the most complex Indian Driving Dataset (IDD) [32] with totally different scene as compared to urban driving Cityscapes in Fig. 6 in Supplementary.
5.2 Plug-and-play application to other domains
@@
| Method | Zero-shot | IoU | F1 | IoU | F1 |
| PROWL | Yes | 8.01 | 14.33 | 62.35 | 76.36 |
| PROWL + CutLER | Yes | 83.29 | 90.38 | 73.30 | 84.08 |
To demonstrate plug-and-play application of PROWL to other domains, we extend our evaluation to rail and maritime ODD scene with respective OOD datasets given in Sec. 4.1. For rail scene, we use in-house created dataset with in-painted OOD objects (see Suppl.). For simplicity, we consider the following predominant classes in RailSem19 [35] for creating prototype bank of ODD classes - train car, platform, rail, fence, person, pole. Similarly, for maritime we refer to MODD [4] and create prototype bank with sky, sea as known ODD classes. Every other object is OOD. Since there exists no benchmark on RailSem19 OOD detection and the MODS [5] uses separate evaluation strategy, we provide evaluation using our zero-shot methods with PROWL. Since CutLER was trained only on unlabelled ImageNet [11], it can easily provide zero-shot inference on any domain, whereas STEGO still needs to train on the datasets of new domain although without labels. Similarly, our supervised methods like Maskomaly [1], RbA [27] relies on Mask2Former which still needs to train with the labels of the respective training data, thus can not be compared.
@@
In Table 3, we show the performance of PROWL vs PROWL with CutLER as a binary segmentation task using fixed default INCS threshold for detection of OOD pixels as provided in Sec. 4.2. We find that PROWL with CutLER performs better as compared to vanilla PROWL in both cases. In Fig. 4, we show corresponding qualitative results. We observe that PROWL is significantly worse in case of Rail Inpainted OOD data while for MODD, PROWL with CutLER performs comparably slightly better than PROWL.This could be attributed to the fact that prototype heatmaps in PROWL provide per-pixel outputs, causing false positives to the background objects like vegetation, buildings that are not in the assumed ODD class list. CutLER provides an advantage here, where the relevant object masks (including OOD) can be filtered based on foreground score while irrelevant background objects can be ignored. However, in MODD dataset, the task is relatively simple as everything other than sea, sky should be detected as OOD or obstacle, thus both methods perform well on this dataset.
5.3 Ablation Study
In Fig. 5, we show the average OOD object size in pixels (a) as well as ablation studies for different hyperparameters used in the experiments with PROWL (b-f). We can see that RA contains much bigger objects compared to RO and FS Static. In (b), we investiagte the influence of different CutLER thresholds for generating optimal number of foreground masks for PROWL with the INCS threshold set at default value of . Optimal mIoU is achieved for CutLER thresholds of , , and for RA, FS Static, and RO datasets respectively. We argue that smaller objects require lower thresholds to be detected reliably but also introduce more detection noise, whereas bigger objects are hard to miss even with a high detector threshold. In (c), we fix the CutLER thresholds given above and find the optimum INCS threshold for best OOD detection based on IoU. We observe FS Static peaks at relatively high threshold of 0.7 as the dataset is quite similar to Cityscapes data, whereas RA and RO dataset peak at , respectively close to the default threshold at . We also compare performance with different DINOv2 model variants as feature extractors in (d). Larger models like ViT-g/14 perform marginally better, however ViT-S/14 seems to offer a better trade-off considering significantly lower inference times. Lastly, we analyze the influence of the prototype samples on the performance in terms of numbers (e) and selection (f). We show that already quite good AUPRC is achieved with as less as prototypes and it starts to saturate from and above. Furthermore, the choice of the specific prototype images seems to have less of an influence as all three distinct sets of samples tested lead to very similar performance. However, further investigation of this property is needed for samples derived from other datasets.